OpenStudy (hanna1):

plz help!! If a 3 month old baby is 20 inches tall, and then at 7 months old, the baby is 26 inches tall, what is the average growth of the baby each month? a. 1.5 in. per month b. 2/3 in. per month c. 6 in. per month d. 6/7 in. per month

OpenStudy (avi13):

26-20/7-4=6/4=1.5 in per month